Consider the follow set of ?=5 (?,?)(x,y) pairs:

  ?x   20 25 26 30 35
  ?y   21 30 35 31 40

NOTE: SXX = 126.8, SXY = 139.6, SYY = 197.2, ?¯x¯ = 27.2, and ?¯y¯ = 31.4.

Compute the sample correlation correct to three decimal places of accuracy.

Correlation = ?=
Answer #1

Answer:
The following table shows the calculations -

x

y

x^2

y^2

xy

20

21

400

441

420

25

30

625

900

750

26

35

676

1225

910

30

31

900

961

930

35

40

1225

1600

1400

Total

136

157

3826

5127

4410

Total number of observations, n = 5

Mean of x, \bar{x} = 136/5 = 27.2 (given)

Mean of y, \bar{y} = 157/5 = 31.4 (given)

Sxx = \sum (x^2) - (n x (\bar{x}^2)) = 3826 – (5 x (27.2^2)) = 126.8 (given)

Syy =  \sum(y^2) - (n x (\bar{y}^2)) = 5127 – (5 x (31.4^2)) = 197.2 (given)

Sxy =  \sumxy - (n x \bar{x} x \bar{y} ) = 4410 – (5 x 27.2 x 31.4) = 139.6 (given)

Therefore, Correlation Coefficient, r = Sxy / (Sxx . Syy)^0.5 = 139.6 / (126.8 x 197.2)^0.5 = 0.883